OpenStudy (anonymous):

What is the equation of the line in standard form that passes through the point (3, –2) and is parallel to the line y = –4x + 1? a) x+4y= 10 b) x-4y= -10 c) 4x-y= -10 d) 4x +y =10

OpenStudy (anonymous):

well, it has to have the slope of -4 since it's parallel. plug in the point. y = -4x + b -2 = -4(3) + b -2 = -12 + b 10 = b so the equation is y = 4x + 10 to get in the right format... y = 4x +10 -10 + y = 4x -10 = 4x - y so it's C.